# EmbAdvisor: Adaptive Cache Management for Sustainable LLM Serving

### v1

@@ 6.1. Methodology
[…]
We define Service Level Objectives (SLOs) for LLM serving performance based on TTFT and TPOT, requiring at least 90 % of requests to meet specified latency targets. For multi-turn conversations, we set TTFT and TPOT thresholds of 2.5 s and 0.2 s for the 70B model, and 0.5 s and 0.15 s for the 8B model. For document reading comprehension—where inputs are longer and latency is less critical, and therefore, we relax the TTFT thresholds to 15 s for the 70B model and 2.5 s for the 8B model. These SLOs are aligned with prior work (Zhong et al., 2024; Agrawal et al., 2024) and benchmarking results from LLMPerf (LLMPerf, 2024).
Comparison points
We evaluate the following system design points.
  • No Cache: Non-caching baseline with vLLM and continuous batching.
  • Full Cache: Use the maximum cache sizes as described in the experiment setup.
  • EmbAdvisor: The carbon-aware caching system in this work. The maximum cache size is the same as Full Cache.
(a) Multi-turn conversation on Llama-3 70B.
(b) Document comprehension on Llama-3 70B.
(c) Multi-turn conversation on Llama-3 8B.
(d) Document comprehension on Llama-3 8B.
Figure 14. SLO attainment and carbon emission under different request rates. Carbon intensity is 124 gCO2e/kWh.
(a) Multi-turn conversation on Llama-3 70B (rate = 1.0 prompt/s).
(b) Document comprehension on Llama-3 70B (rate = 0.2 prompt/s).
(c) Multi-turn conversation on Llama-3 8B (rate = 3.5 prompt/s).
(d) Document comprehension on Llama-3 8B (rate = 2.0 prompt/s).
Figure 15. SLO attainment and carbon emission under different carbon intensities (CIs).
6.2. SLO Attainment and Carbon Emission
We evaluate SLO attainment and carbon emissions of EmbAdvisor under different request rates and carbon intensities.
6.2.1. Carbon and Performance vs. Request Rate
We first conduct an experiment with the variable request rates. For each task and model, we start from a lower rate to a higher rate, beyond which the SLOs can no longer be satisfied. We take the average carbon intensity (CI) in 2024 of ES grid, which is 124 gCO2e/kWh. Figures 14(b), 14(a), 14(d) and 14(c) compares the performance and carbon emissions of EmbAdvisor and two baselines, No Cache and Full Cache. Performance-wise, EmbAdvisor and Full Cache meets SLO all the time (i.e., 90 % of requests meet both TTFT and TPOT SLOs), while No Cache sometimes breaks SLOs as the request rate increases (indicated by the red “✘” sign). Under low request rates, EmbAdvisor demonstrates better carbon emissions than other scenarios, while ensuring 90 % of requests meet SLOs. Compared to Full Cache, in the two tasks, the 70B model achieves up to 11.1 % and 11.0 % carbon reductions, respectively; the 8B model achieves up to 8.2 % and 2.5 % carbon reductions, respectively. The benefit is mainly because EmbAdvisor allocates smaller cache sizes, reducing the embodied carbon of SSD.
As the request rate becomes higher, EmbAdvisor tends to choose larger cache sizes or even the full cache to satisfy SLOs. In such cases, No Cache fails to meet SLO. Specifically, the 8B model fails to meet SLO in document reading comprehension when request rate is 2.5 prompt/s even with Full Cache. Overall, compared to Full Cache, EmbAdvisor reduces the carbon emissions by 6.4 % and 4.2 % on average in 70B and 8B models, respectively. while meeting SLOs.
6.2.2. Carbon and Performance vs. Carbon Intensity
We also evaluate the balance between carbon and performance under different carbon intensities (CIs). We take the average carbon intensities in 2024 of the 4 grids in Figure 2 and use the middle request rates in Section 6.2.1. In reality, carbon intensity can change significantly not only across regions but also within a day, as discussed in Section 3.2.2. Therefore, adaptivity to carbon intensity is critical. Figures 15(b), 15(a), 15(d) and 15(c) demonstrate the results. When CI is low (FR grid), No Cache has the lowest carbon emission as embodied carbon has a higher weight, but fails to meet SLOs most of the time. In comparison, EmbAdvisor still ensures 90 % SLO attainment. Compared to Full Cache, EmbAdvisor reduces carbon emissions up to 31.2 %, 28.6 % of the 70B model in the two tasks, respectively, and 13.6 % of the 8B model in the multi-turn conversation task. EmbAdvisor reduces more carbon emissions for the 70B model, primarily due to higher storage reduction — at least 12 TB smaller cache size when the CI is 33 gCO2e/kWh. The document reading comprehension task is more intensive and thus, when running 8B mode, EmbAdvisor always uses the maximum cache size, without distinction from the Full cache configuration.
As CI increases, the operational carbon dominates. Therefore, larger cache sizes achieve better performance and yields lower total carbon despite higher embodied carbon. The SLO attainment remains constant under No Cache and Full Cache configurations, as they are unaware of the carbon intensity. EmbAdvisor demonstrates SLO attainment changes as shown in Figure 15(b) because of the adaptive cache size configuration. Compared to the 8B model, the 70B model requires more storage to achieve the best performance. Therefore, EmbAdvisor reduces more storage in the 70B model, achieving more embodied carbon reduction. Overall, compared to Full Cache, EmbAdvisor reduces carbon by 9.5% and 3.7% for the 70B and 8B models, respectively.
6.3. Ablation Study
(a) Multi-turn conversation.
(b) Document comprehension.
Figure 16. Ablation study on adaptive caching.
6.3.1. Adaptive Caching Analysis
We first conduct an experiment that integrates EmbAdvisor’s adaptive caching technique into LMCache (LMCache Team, 2025b) while using its original LRU policy (i.e., “LRU+Optimal”). Figure 16 shows the carbon reduction under different request rates. We follow the carbon intensity in Section 6.2.1. Compared to Full Cache, EmbAdvisor’s adaptive caching strategy can bring up to 10.3 % carbon benefits in multi-turn conversation and 6.6 % reduction in document reading comprehension. As the request rate increases, the benefit reduces, as EmbAdvisor adapts to larger cache sizes to meet SLOs.
6.3.2. Replacement Policy Comparison
We then compare the hit rates of different replace policies, as listed below.
  • First In, First Out (FIFO): Replace blocks in the order in which they were added as a FIFO queue.
  • Least Recently Used (LRU): Replace the least recently used cache block. It is the default policy in LMCache (LMCache Team, 2025b).
  • Carbon-saving-aware (CSA): The replacement policy of EmbAdvisor that replaces the entry that brings the least carbon emission savings (details in Section 5.5).
We evaluate cache sizes ranging from 1 to 16 TB with the Llama-3 70B model. We define cache hit rate as the number of tokens reused from the cache over the total number of input tokens, as shown in Table 3. In the vast majority of cases, our replacement policy CSA outperforms other replacement policies. Although LRU indicates similar performance as CSA with 16TB cache size, CSA demonstrates up to 9 % higher hit rate than LRU with smaller cache sizes.
Table 3. Hit rate comparison (Llama-3 70B).
Cache Size (TB) ShareGPT TriviaQA
@@ 6.3.2. Replacement Policy Comparison
CSA
1 0.05 0.05 0.08 0.10 0.11 0.19
2 0.11 0.12 0.17 0.16 0.19 0.25
4 0.19 0.21 0.28 0.23 0.26 0.35
8 0.34 0.40 0.47 0.37 0.38 0.44
12 0.38 0.55 0.59 0.41 0.45 0.48
16 0.52 0.69 0.71 0.47 0.52 0.52
[…]
6.4. Sensitivity Study
6.4.1. Document Reading Comprehension Skewness
[…]
6.4.2. Sensitivity Study on SSD Lifespan
@@
(b) Document comprehension.
Figure 18. Variable SSD lifespan (Llama-3 70B).
Prior studies have found that the lifespans of SSDs vary by manufacturer and usage. In this experiment, we study the carbon emission savings from EmbAdvisor by varying the SSD lifetime from 3 to 7 years (McAllister et al., 2024a; Lyu et al., 2023b; Acun et al., 2023; Lyu et al., 2023a). We set the request rate as 1.5 prompt/s for the multi-turn conversation task and 0.2 prompt/s for the document comprehension task. We use the carbon intensity of the ES grid, same as Section 6.2.1. Figure 18 indicates that a shorter, 3-year SSD lifetime leads to more carbon savings from EmbAdvisor, with up to 11.9 %. With the extension of SSD’s lifetime, the amortized embodied carbon declines. Therefore, SSD takes a smaller proportion of the total carbon emissions, resulting in less carbon savings from resizing the cache.
6.4.3. Sensitivity Study on Storage Embodied Carbon
@@
(b) Document comprehension.
Figure 19. Variable SSD embodied carbon (Llama-3 70B).
@@
In this work, we model SSD embodied carbon using ACT (Gupta et al., 2022). SSD embodied carbon varies by factors such as the manufacturing technology, fab location, and account scope, and recent works have reported higher embodied carbon of SSDs (Tannu and Nair, 2023; Bhagavathula et al., 2024). Therefore, we further evaluate a range of embodied carbon emissions of SSDs, from our assumed 30 kgCO2e/TB to a higher value of 90 kgCO2e/TB. We evaluate two datasets with request rates of 1.5 and 0.2 prompt/s, respectively. We also follow the same CI as Section 6.2.1. Figure 19 demonstrates that with embodied carbon of SSD increasing, the carbon benefits from EmbAdvisor also increase as embodied carbon takes a larger part of the total emissions. EmbAdvisor can reduce more embodied carbon by shrinking the cache size. Overall, EmbAdvisor saves up to 25 % carbon if the SSD’s embodied carbon is 90 kgCO2e/TB.
@@ 7. Related Work
Caching for LLM. Storing KV caches among different requests is a common optimization (Gao et al., 2024; Liu et al., 2024a; Gao et al., 2025). For example, CachedAttention (Gao et al., 2024) proposes that the KV cache can be reused among multi-turn conversations, by using TBs of caches. CacheGen (Liu et al., 2024a) reduces the network overhead by compressing the transferred KV caches when the SSD is deployed remotely. HCache (Gao et al., 2025) restores historical LLM states with intermediate activations and manages the storage at chunk-wise to alleviate the overhead from I/O or recomputing. Instead of the precise matching in tokens, prior works also propose semantic caching. For example, GPTCache (Bang, 2023) retrieves the cache by its semantics and returns the same response as the retrieved cache. Unlike these studies that only optimizes for performance, EmbAdvisor’s achieves a tradeoff between the performance and carbon emissions. The existing caching solutions can also be integrated into EmbAdvisor framework to reduce their total carbon emissions.
Carbon mitigation for LLM. Prior works have proposed various solutions to alleviate carbon emissions of LLMs (Stojkovic et al., 2025; Faiz et al., 2024; Shi et al., 2024; Samsi et al., 2023). For example, DynamoLLM (Stojkovic et al., 2025) schedules and manages requests on various GPUs under diverse frequency settings to minimize operational carbon while satisfying SLOs. LLMCarbon (Faiz et al., 2024) builds a carbon model to predict the emissions due to LLM training and inference. GreenLLM (Shi et al., 2024) notices the heterogeneous demand in LLM and proposes to use different types of GPUs to lower down both embodied and operational carbon. From Words to Watts (Samsi et al., 2023) benchmarks LLM on various settings to show the energy and performance efficiency. EcoServe (Li et al., 2025) analyzes the proportion of the operational and embodied carbon, and schedules the LLM applications with resource provision. These works provide substantial carbon emission savings from the computation side but overlook the importance of the embodied carbon from storage, which is EmbAdvisor’s main angle.
@@ 8. Conclusions
The wide use of large language models (LLMs) leads to high carbon emissions. While most LLM sustainability studies focus on compute-related emissions, we find that storage for saving the KV cache of context in previous prompts is another significant contributor. Although caching improves performance and reduces operational emissions, it also introduces substantial embodied carbon due to the high-speed, high-capacity SSDs. To address this tradeoff, we introduce EmbAdvisor, a carbon-aware caching framework that dynamically profiles LLM tasks and uses an ILP-based optimizer to determine the optimal cache size, balancing performance and carbon efficiency while meeting SLOs.

### v2

@@ 6.3.2. Replacement Policy Comparison
LCS
FIFO
LRU
LCS
1 0.05 0.05 0.08 0.05 0.05 0.06 0.10 0.11 0.19
2 0.11 0.12 0.17 0.08 0.08 0.09 0.16 0.19 0.25
4 0.19 0.21 0.28 0.12 0.12 0.14 0.23 0.26 0.35
8 0.34 0.40 0.47 0.19 0.20 0.22 0.37 0.38 0.44
16 0.52 0.69 0.71 0.31 0.32 0.33 0.47 0.52 0.52
Figure 19. Constraint solver execution time for every cache resize.
6.4. Constraint Solver Overhead
We evaluate the execution time of the constraint solver when making each decision, as shown in Figure 19. The average latency of making a cache resizing decision is as low as 7.03 s. This is a low overhead given that these serving workloads are long-running.
@@ 6.4. Constraint Solver Overhead
(b) Document comp. ($\alpha=0.4$).
(c) Document comp. ($\alpha=0.7$).
Figure 20. Impact of prediction and profile inaccuracies (Llama-3 70B).
6.5. Predictors and Profiler Errors
GreenCache decisions can be affected by three sources of errors, CI predictor, load predictor, and profiler as discussed section 5.4. We first calculate the Mean Average Percentage Error (MAPE) of both predictors by comparing the predictions with the groundtruth. The load predictor has an MAPE of 4.3 %, and the CI predictor has MAPE values of 12.7 %, 15.3 %, 11.3 %, and 6.8 % for FR, FI, ES, and CISO grids, respectively. Then, we calculate the distribution differences between the profiling data and the evaluation data. In multi-turn conversation and document comprehension tasks, the median context length difference is 5.78 % and 1.13 %, respectively.
We further evaluate the impact on carbon emission savings by comparing the carbon emissions of GreenCache under errors with an ideal scenario that uses the groundtruth. Figure 20 shows the reduction of carbon emission savings due to errors. To make small values visible, we plot the values on a log scale. CI prediction errors only reduce 0.0064 % of carbon savings on average compared to the ideal scenario. When load prediction errors are included, the reduction is increased to an average of 0.20 %. Although the load predictor has lower MAPEs than CI, it directly affects the choice of cache size and thus has a higher impact on carbon emissions. Finally, the errors from the profiler increase the carbon savings reduction to an average of 0.79 %. The profiler errors have a relatively higher impact in low-CI grids (FR and FI), as embodied carbon has a higher weight in these grids. We conclude that these errors have an overall low impact on carbon emission savings.
6.6. Sensitivity Studies
In this section, we vary the decision-making frequency, SSD lifespan, and SSD embodied carbon to analyze their impact on carbon emissions.
@@ 6.6. Sensitivity Studies
(b) Document comp. ($\alpha=0.4$).
(c) Document comp. ($\alpha=0.7$).
Figure 21. Impact of variable cache resizing intervals (Llama-3 70B). A higher value indicates more savings.
6.6.1. Cache Resizing Interval
We evaluate the carbon emission savings of Llama-3 70B over full cache under variable resizing intervals, in addition to the default 1-hour interval. Figure 21 shows the relative difference compared to the default 1-hour interval. We also adjust the CI and rate predictors to match the granularity. Because the CI data has a minimum granularity of 1 hour due to the dataset, we use the fixed hourly CI value instead when the interval is within 1 hour. GreenCache sets a sufficiently large cache size during the whole interval to ensure the SLO attainment goal. Therefore, the carbon emission savings are significantly reduced under longer intervals across all grids. Among the tasks, the difference is more significant in the document comprehension that requires larger caches, especially under a lower skewness of $\alpha=0.4$.
6.6.2. SSD Lifespan
(a) Multi-turn conversation.
(b) Document comp. ($\alpha=0.4$).
(c) Document comp. ($\alpha=0.7$).
Figure 22. Variable SSD lifespan (Llama-3 70B, ES grid).
(a) Multi-turn conversation.
(b) Document comp. ($\alpha=0.4$).
(c) Document comp. ($\alpha=0.7$).
Figure 23. Variable SSD embodied carbon (Llama-3 70B, ES grid).
Prior studies have found that the lifespans of SSDs vary by manufacturer and usage. In this experiment, we study the carbon emission savings from GreenCache by varying the SSD lifetime from 3 to 7 years (McAllister et al., 2024a; Lyu et al., 2023b; Acun et al., 2023; Lyu et al., 2023a). We use a fixed rate of 1.5 prompts/s for the multi-turn conversation task and 0.2 prompts/s for document comprehension. Like the ablation studies, we also use the average carbon intensity of the ES grid. Figure 22 indicates that a shorter, 3-year SSD lifetime leads to more carbon savings from GreenCache, with up to 11.9 %. Extending SSD lifetime reduces amortized embodied carbon, lowering the savings from resizing the cache.
6.6.3. SSD Embodied Carbon
@@
In this work, we model SSD embodied carbon using ACT (Gupta et al., 2022). SSD embodied carbon varies by factors such as the manufacturing technology, fab location, and account scope. Recent studies have reported higher embodied carbon of SSDs (Tannu and Nair, 2023; Bhagavathula et al., 2024). Therefore, we evaluate a range of embodied carbon emissions of SSDs, from 30 to 90kgCO2e/TB. Same as the previous sensitivity study, we evaluate the two tasks with fixed rates of 1.5 and 0.2 prompts/s, respectively. We also follow the CI of the ES grid. Figure 23 demonstrates that with the embodied carbon of SSD increasing, the carbon benefits from GreenCache also increase as embodied carbon has a higher weight in the total emissions. GreenCache can reduce more embodied carbon by shrinking the cache size. Overall, GreenCache saves up to 25 % carbon if the SSD’s embodied carbon is 90 kgCO2e/TB.
7. Discussion
In this section, we discuss the assumptions of embodied carbon accounting, potential improvement from more advanced load and CI predictors, and implications on MoE models.
Embodied Carbon Accounting. Consistent with prior work (Han et al., 2025; Wang et al., 2025; Li et al., 2025; Shi et al., 2024, 2025), we consider both embodied and operational carbon emissions in cloud environments, emphasizing that embodied carbon is a critical component that should be included. Different from the assumptions in Sunk Carbon Fallacy (Bashir et al., 2024), which treat unselected resources as idle and their embodied carbon as a fixed sunk cost, this work assumes a cloud scenario in which unused resources are promptly utilized by other workloads. Consequently, we model embodied carbon as continuously amortized across workloads over time and explicitly account for it in resource provisioning and management. In particular, for storage that this work focuses on, GreenCache dynamically resizes storage, and thus the embodied carbon of storage (i.e., SSD) is also attributed based on the actual storage provision.
Load Prediction. We design a lightweight load predictor, as discussed in Section 5.3. The main limitation of the request rate study is the lack of real-world LLM serving traces, as the publicly available Azure trace only covers 1 week. With more trace data, it is possible to achieve higher prediction accuracy, like prior works in other cloud computing domains (Zhang et al., 2024; Li et al., 2011; Zhang et al., 2025), which can lead to even better carbon emission savings.
Carbon Intensity Forecast. GreenCache uses a recent, state-of-the-art CI prediction model, EnsembleCI (Yan et al., 2025). CI prediction is a well-known problem in sustainable computing (Maji et al., 2023, 2022; Yan et al., 2025). Like the load predictor, the CI predictor is an easily replaceable module in the GreenCache framework. With a more accurate prediction model, we expect more accurate cache allocation.
Implications on MoE models. Mixture-of-Expert (MoE) models are widely deployed (Dai et al., 2024; Team, 2024). Compared to dense LLMs, MoE models incorporate multiple FFNs as “experts” and activate only a subset of them, leading to lower computation overhead. On the other hand, the KV cache storage is not reduced. Therefore, MoE models lower operational carbon emissions but amplify the significance of embodied carbon, making the optimizations provided by GreenCache more impactful.
@@ 8. Related Work
In this section, we discuss caching for LLM systems and other carbon mitigation strategies.
Caching for LLM. Storing KV caches among different requests is a common optimization (Gao et al., 2024; Liu et al., 2024b; Gao et al., 2025). For example, CachedAttention (Gao et al., 2024) proposes to reuse the KV cache among multi-turn conversations with TBs of caches. CacheGen (Liu et al., 2024b) reduces the network overhead by compressing the transferred KV caches when the SSD is deployed remotely. HCache (Gao et al., 2025) restores historical LLM states with intermediate activations and manages the chunk-wise storage to alleviate the overhead from I/O or recomputing. Instead of the precise matching in tokens, prior works also propose semantic caching. For example, GPTCache (Bang, 2023) retrieves the cache by its semantics and returns corresponding responses. Unlike these studies that only target performance, GreenCache achieves a tradeoff between performance and carbon emissions. The existing caching solutions can also be integrated into GreenCache framework to reduce their total carbon emissions.
Carbon mitigation for LLM. Prior works have proposed various solutions to alleviate carbon emissions of LLMs (Stojkovic et al., 2025; Faiz et al., 2024; Shi et al., 2024, 2025; Samsi et al., 2023). For example, DynamoLLM (Stojkovic et al., 2025) schedules and manages requests on various GPUs under diverse frequency settings to minimize operational carbon while satisfying SLOs. LLMCarbon (Faiz et al., 2024) builds a carbon model to predict the emissions due to LLM training and inference. GreenLLM (Shi et al., 2024, 2025) notices the heterogeneous demand in LLM and proposes to use different types of GPUs to lower both embodied and operational carbon. From Words to Watts (Samsi et al., 2023) benchmarks LLM on various settings to show the energy and performance efficiency. EcoServe (Li et al., 2025) analyzes the proportion of the operational and embodied carbon, and schedules the LLM applications with resource provision. These works provide substantial carbon emission savings from the computation side but overlook the importance of the embodied carbon from storage, which is GreenCache’s main angle.
@@ 9. Conclusions
The wide use of large language models (LLMs) leads to high carbon emissions. While most LLM sustainability studies focus on compute-related emissions, we find that storage for saving the historical KV cache is another significant contributor. Although caching improves performance and reduces operational emissions, it also introduces substantial embodied carbon due to the high-speed, high-capacity SSDs. To address this tradeoff, we introduce GreenCache, a carbon-aware caching framework that dynamically profiles LLM tasks and uses an ILP-based optimizer to reconfigure cache size, balancing performance and carbon efficiency while meeting the SLO attainment goal.
Acknowledgement
We thank the anonymous reviewers and the shepherd for their valuable feedback on this work, and Shuncheng Jie and Dengwang Tang for proofreading. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) RGPIN-2023-03478. We also acknowledge the support of the U.S. National Science Foundation (NSF) CCF-2413870.
Appendix A ILP Complexity
Theorem 0 (NP-hardness).
The GreenCache optimization problem in Eq. (5.4.1) is NP-hard, even in a restricted setting where each time step only allows a binary cache decision (off/on), and the SLO requirement is a global ratio constraint: at least a configurable fraction $\rho$ of all requests over the horizon satisfy the TTFT threshold and at least the same fraction $\rho$ satisfy the TPOT threshold (where $\rho$ is part of the input).
Proof.
We prove NP-hardness by a polynomial-time reduction from the NP-complete 0–1 Knapsack decision problem. A Knapsack instance consists of $m$ items, where item $k$ has weight $w_{k}\in\mathbb{Z}_{>0}$ and value $v_{k}\in\mathbb{Z}_{>0}$, a weight budget $W$, and a target value $V$. The question is whether there exists a subset $K\subseteq\{1,\dots,m\}$ such that $\sum_{k\in K}w_{k}\leq W$ and $\sum_{k\in K}v_{k}\geq V$.
Decision version of GreenCache. Consider the decision form: given a carbon budget $C$, is there a cache plan such that total carbon $\leq C$ and the (global-$\rho$) SLO constraints hold?
Construction. Given a Knapsack instance, we construct a restricted GreenCache instance as follows.
  • Set $T=m$, which maps the number of items ($m$) in the Knapsack problem directly to the total number of time steps ($T$) in the GreenCache instance. Map item $k$ to time step $t=k$. Restrict the cache decision at each step to $S_{k}\in\{0,1\}$ (cache off/on).
  • The Knapsack target value $V$ maps to the minimum number of requests that must satisfy the SLO constraints in the GreenCache instance.
  • Set the request volume at time step $k$ to be $\lambda_{k}:=v_{k}$. Let $\Lambda:=\sum_{k=1}^{m}\lambda_{k}=\sum_{k=1}^{m}v_{k}$. If $V>\Lambda$, the knapsack instance is trivially infeasible; in this case we output any GreenCache instance that is trivially infeasible (e.g., by setting an impossible SLO), and the reduction remains correct. Hence, assume $V\leq\Lambda$.
  • Fix the workload SLO thresholds $\theta_{\text{TTFT}}$ and $\theta_{\text{TPOT}}$. For each step $k$, let $q^{\text{TTFT}}_{k}$ (resp. $q^{\text{TPOT}}_{k}$) denote the number of requests at step $k$ whose TTFT (resp. TPOT) meets the corresponding threshold. We enforce that at least a fraction $\rho$ of all requests satisfy each threshold:
    $\sum_{k=1}^{m}q^{\text{TTFT}}_{k}\geq\rho\Lambda,\qquad\sum_{k=1}^{m}q^{\text{TPOT}}_{k}\geq\rho\Lambda.$
    We set the (input) ratio parameter to
    $\rho:=\frac{V}{\Lambda}\in(0,1].$
    Hence $\rho\Lambda=V$, and the global-$\rho$ constraints become $\sum_{k}q^{\text{TTFT}}_{k}\geq V$ and $\sum_{k}q^{\text{TPOT}}_{k}\geq V$.
  • We choose the instance’s latency functions/coefficients so that, for each step $k$:
    1. (1)
      If $S_{k}=1$, then all $\lambda_{k}$ requests have TTFT and TPOT below the corresponding thresholds, i.e., $q^{\text{TTFT}}_{k}=q^{\text{TPOT}}_{k}=\lambda_{k}$.
    2. (2)
      If $S_{k}=0$, then all $\lambda_{k}$ requests violate at least one of the two thresholds, so $q^{\text{TTFT}}_{k}=q^{\text{TPOT}}_{k}=0$.
    Therefore, the number of requests satisfying each threshold at step $k$ is $q^{\text{TTFT}}_{k}=q^{\text{TPOT}}_{k}=\lambda_{k}S_{k}$.
  • Set the incremental carbon at time step $k$ to be exactly $w_{k}\cdot S_{k}$, and set the carbon budget to $C:=W$. Any carbon terms that are independent of the cache decisions (e.g., constant embodied carbons) are in the budget $C$ without affecting NP-hardness.
Correctness. By construction, for every $k$ we have $q^{\text{TTFT}}_{k}=q^{\text{TPOT}}_{k}=\lambda_{k}S_{k}$. Thus, the global SLO constraints become
$\sum_{k=1}^{m}\lambda_{k}S_{k}\geq V\quad\Longleftrightarrow\quad\sum_{k=1}^{m}v_{k}S_{k}\geq V.$
Meanwhile, the carbon budget constraint is exactly
$\sum_{k=1}^{m}w_{k}S_{k}\leq W.$
Therefore, there exists a feasible solution to the constructed GreenCache decision instance iff there exists a feasible knapsack subset.
Polynomial-time reduction. The construction sets $T=m$ and uses only $\{(w_{k},v_{k})\}_{k=1}^{m}$ and simple arithmetic. The rational parameter $\rho=V/\Lambda$ has polynomial bit-length (it can be represented by the integer pair $(V,\Lambda)$). Hence the reduction runs in polynomial time.
Therefore, the GreenCache problem is NP-hard even in this restricted special case; consequently, the general optimization problem in Section 5.4.1 is NP-hard. ∎